# EasyHOI: Unleashing the Power of Large Models for Reconstructing Hand-Object Interactions in the Wild

### v3

@@ 3.3 Initial Reconstruction of Hand and Object
Object Reconstruction. Due to the occlusion caused by hand, the interacted object is often visually incomplete in the input image and therefore could only produce distorted 3D geometry of the object. To address this problem, as the example shown in Fig. 3, our reconstruction pipeline first employs a diffusion model [70, 43] to recover the complete object appearance by removing occlusions. As a by-product, we could obtain the segmentation mask of the impainted object with complete appearance, denoted as $\hat{M}_{o}$.
@@ 3.4 Hand-Object Interaction Optimization
1) Camera System Setup. The problem of aligning the two systems of hand and object can be simplified by using the reconstructed object’s coordinate frame as the global reference, and then estimating the camera parameters of the input image defined in the object coordinate system. Using a differentiable renderer $\Psi$, we would like to obtain the optimal camera parameters ($K$, $[R|t]$) by minimizing the soft IoU loss between rendered object mask $M_{o}^{r}=\Psi(\Omega_{o},K,R,t)$ and ground truth mask $\hat{M}_{o}$ (the segmentation mask of the impainted object with complete appearance), as shown in Eq. 1.
@@ 4 Experiments
Table 1: Quantitative evaluation for HOI reconstruction. Since AlignSDF and gSDF were trained on DexYCB, we exclude their DexYCB results from our zero-shot comparisons. The metrics F5 and F10 measure the F score of points from reconstructed object within 5mm and 10mm of the GT object, respectively. The metric $C.D.$ denotes the Chamfer Distance between reconstructed object and GT object, $S.D.$ denotes Simulation Displacement(in $cm$) and $I.V.$ represents Intersection Volume(in $cm^{3}$).
@@ 4.2 Datasets and Baselines
Camera HOI Contact Hand Parameter Sim.Disp.$(cm)\downarrow$ Int.Vol.$(cm^{3})\downarrow$ Sim.Disp.$(cm)\downarrow$ Int.Vol.$(cm^{3})\downarrow$ Sim.Disp.$(cm)\downarrow$ Int.Vol.$(cm^{3})\downarrow$
@@ 4.4 Comparison Results
Quantitative Results. We evaluate our method’s performance in terms of object reconstruction accuracy and grasping quality. To ensure fair comparison, we employ HaMeR for hand pose detection when evaluating all baseline methods. As shown in Tab. 1, our method achieves the lowest Chamfer distance and highest F-scores. These metrics demonstrate our method’s capability to accurately reconstruct objects from single-view images. In terms of hand-object interaction quality, Tab. 1 shows that our method outperforms competing approaches, consistently achieving the lowest simulation displacement and smallest intersection volume across all datasets. These results underscore the effectiveness of our hand-object interaction modeling.
@@ 4.5 Ablation Study and Discussions
Full losses 3.08 2.92 4.11 4.77
w/o $\mathcal{L}_{\text{hand-mask}}$ 3.21 3.69 4.28 4.69
w/o $\mathcal{L}_{\text{penetr}}$ 2.95 3.16 9.62 10.13
w/o $\mathcal{L}_{\text{reg}}$ 3.24 3.73 4.47 4.82
w/o $\mathcal{L}_{\text{contact}}$ 3.94 3.81 4.26 4.79
@@
Ablation of Loss Terms in Hand Parameter Refinement. To further examine Hand Parameter Refinement, we conducted an ablation study on the loss terms, with results shown in  Fig. 7 and Tab. 3. Both qualitative and quantitative evaluations offer insights into the significance of each loss term. Although $\mathcal{L}_{\text{hand-mask}}$ was optimized in a prior step, its removal still increases simulation displacement and introduces deviations from the input image. Additionally, removing $\mathcal{L}_{\text{penetr}}$ results in more intersections, highlighting its essential role in maintaining physical realism. Without $\mathcal{L}_{\text{reg}}$, we observe twisted fingers, higher simulation displacement, and larger intersection volumes, which reduce grasp fidelity. Lastly, omitting $\mathcal{L}_{\text{contact}}$ positions the fingertips farther from the object surface and increases simulation displacement, underscoring its critical role in preserving grasp stability. These findings confirm that each loss term is crucial for achieving physically plausible and stable hand-object interactions.
@@ 6 Acknowledgements
This supplementary material provides additional details on our method and results that complement the main paper. Sec. 7 details the segmentation and reconstruction of the object. Sec. 8 elaborates on the HOI contact alignment stage of the HOI optimization process. Finally, Sec. 9 presents further comparative analyses on public datasets, demonstrating the robustness and versatility of our method.
@@ 9.1 Qualitative Comparison Results
The Selection of Large Reconstruction Models.

### v4

@@ 3.3 Initial Reconstruction of Hand and Object
Object Reconstruction. Due to the occlusion caused by hand, the interacted object is often visually incomplete in the input image and therefore could only produce distorted 3D geometry of the object. To address this problem, as the example shown in Fig. 3, our reconstruction pipeline first employs a diffusion model [71, 43] to recover the complete object appearance by removing occlusions. As a by-product, we could obtain the segmentation mask of the inpainted object with complete appearance, denoted as $\hat{M}_{o}$.
@@ 3.4 Hand-Object Interaction Optimization
1) Camera System Setup. The problem of aligning the two systems of hand and object can be simplified by using the reconstructed object’s coordinate frame as the global reference, and then estimating the camera parameters of the input image defined in the object coordinate system. Using a differentiable renderer $\Psi$, we would like to obtain the optimal camera parameters ($K$, $[R|t]$) by minimizing the soft IoU loss between rendered object mask $M_{o}^{r}=\Psi(\Omega_{o},K,R,t)$ and ground truth mask $\hat{M}_{o}$ (the segmentation mask of the inpainted object with complete appearance), as shown in Eq. 1.
@@ 4 Experiments
Table 1: Quantitative evaluation for object quality in HOI reconstruction. Since AlignSDF and gSDF were trained on DexYCB, we exclude their DexYCB results from our zero-shot comparisons. The metrics F5 and F10 measure the F score of points from reconstructed object within 5mm and 10mm of the GT object, respectively. The metric $C.D.$ denotes the Chamfer Distance between reconstructed object and GT object, $S.D.$ denotes Simulation Displacement(in $cm$) and $I.V.$ represents Intersection Volume(in $cm^{3}$).
Arctic OakInk
[…]
IHOI 1.36 1.57 1.36 1.32
AlignSDF 1.38 1.36 1.25 1.26
gSDF 1.34 1.37 1.27 1.14
MOHO 1.38 1.35 1.27 1.33
HaMeR 1.14 1.05 1.06 1.03
Ours (Pred-object) 1.48 1.36 1.19 1.22
Ours (GT-object) 1.12 0.95 0.86 0.77
Table 2: Hand pose accuracy in HOI reconstruction is evaluated using MPVPE (Mean Per Vertex Position Error) and MPJPE (Mean Per Joint Position Error), both measured in centimeters (cm).
@@ 4.2 Datasets and Baselines
Camera HOI Contact Hand Parameter S.D.$\downarrow$ I.V.$\downarrow$ S.D.$\downarrow$ I.V.$\downarrow$ S.D.$\downarrow$ I.V.$\downarrow$
@@ 4.4 Comparison Results
Quantitative Results. We evaluate our method’s performance in terms of object reconstruction accuracy, grasping quality and hand accuracy. To ensure fair comparison, we employ HaMeR for hand pose detection when evaluating all baseline methods. As shown in Tab. 1, our method achieves the lowest Chamfer distance and highest F-scores. These metrics demonstrate our method’s capability to accurately reconstruct objects from single-view images. In terms of hand-object interaction quality, Tab. 1 shows that our method outperforms competing approaches, consistently achieving the lowest simulation displacement and smallest intersection volume across all datasets. These results underscore the effectiveness of our hand-object interaction modeling. In Tab. 2, we report the reconstruction errors for hand vertices and joints. Since the object geometry recovered by LRM is not entirely accurate, optimizing for physically plausible HOI results with such imperfect object geometry inevitably compromises hand accuracy. Consequently, our final hand results are slightly worse than HaMER’s. However, with ground truth object geometry, our proposed HOI optimization can further refine the initial hand results produced by HaMER, demonstrating the effectiveness of our method.
@@ 4.5 Ablation Study and Discussions
Full losses 3.08 2.92 4.11 4.77 1.19 1.22
w/o $\mathcal{L}_{\text{hand-mask}}$ 3.21 3.69 4.28 4.69 1.73 1.68
w/o $\mathcal{L}_{\text{penetr}}$ 2.95 3.16 9.62 10.13 1.35 1.33
w/o $\mathcal{L}_{\text{reg}}$ 3.24 3.73 4.47 4.82 4.57 4.41
w/o $\mathcal{L}_{\text{contact}}$ 3.94 3.81 4.26 4.79 1.30 1.27
@@
Ablation of Loss Terms in Hand Parameter Refinement. To further examine Hand Parameter Refinement, we conducted an ablation study on the loss terms, with results shown in  Fig. 7 and Tab. 4. Both qualitative and quantitative evaluations offer insights into the significance of each loss term. Although $\mathcal{L}_{\text{hand-mask}}$ was optimized in a prior step, its removal still increases simulation displacement and introduces deviations of hand shape from the input image. Additionally, removing $\mathcal{L}_{\text{penetr}}$ results in more intersections, highlighting its essential role in maintaining physical realism. Without $\mathcal{L}_{\text{reg}}$, we observe twisted fingers, higher simulation displacement, and larger intersection volumes, which reduce grasp fidelity. Lastly, omitting $\mathcal{L}_{\text{contact}}$ positions the fingertips farther from the object surface and increases simulation displacement, underscoring its critical role in preserving grasp stability. These findings confirm that each loss term is crucial for achieving physically plausible and stable hand-object interactions.
@@ 6 Acknowledgements
This supplementary material provides additional details on our method and results that complement the main paper. Sec. 7 details the segmentation and reconstruction of the object. Sec. 8 elaborates on the HOI contact alignment stage of the HOI optimization process. Finally, Sec. 9 presents further experiments and analyses, demonstrating the robustness and versatility of our method.
@@ 9 Experiments
9.1 Running Time Analysis
We measured the processing time for 65 images at each stage and calculated the average runtime. The initial reconstruction stage required an average of 58.97 seconds, with the breakdown as follows: HOI Reasoning (11.68 seconds), Hand Reconstruction (5.54 seconds), Object Inpainting (15.55 seconds), and Object Reconstruction (26.20 seconds). For the optimization stage, the average runtime was 57.03 seconds, consisting of Camera Setup (4.41 seconds), Contact Alignment (30.51 seconds), and Hand Refinement (22.11 seconds).
9.2 Ablation on Large Models
The Selection of HOI Reasoning Models.
In addition to the vision-language model, we examine the Hand-Object detector (HODet)[53] for segmenting hands and objects in input images. To compare HOI reasoning performance, we evaluate LISA against HODet on the Arctic dataset. As HODet predicts only object bounding boxes, we employ SAM to generate object segmentations based on these predictions. LISA outperforms HODet, achieving an average IoU of 0.74 compared to 0.61. A visual comparison in Figure 12 shows that HODet frequently misidentifies objects and detects extraneous elements.
[…]
Figure 12: Visualization of the ablation on HOI reasoning large model.
The Selection of Large Reconstruction Models.
To assess the impact of reconstruction quality, we evaluate the optimization results using InstantMesh reconstruction, Tripo3D reconstruction, and ground truth (GT) meshes on Oakink dataset. As shown in the table below, lower-quality reconstructions (indicated by higher Chamfer Distance (C.D.) values) result in poorer HOI performance.
Object quality HOI results final hand results
[…]
GT 1.00 1.00 0.00 1.92 2.44 0.86 0.77
Tripo3D 0.280 0.503 0.842 2.76 4.05 1.15 1.21
InstantMesh 0.247 0.445 1.035 3.08 4.11 1.19 1.22
Table 5: The impact of object reconstruction quality on HOI optimization performance.
@@ 9.3 Qualitative Comparison Results
Comparative results on In-th-wild Images
As depicted in Figure 13, we showcase qualitative comparisons with IHOI, AlignSDF, gSDF, MOHO, and our method on in-the-wild images. Our approach excels in accurately reconstructing intricate object geometries and details.
Figure 13: Qualitative Comparison of IHOI, AlignSDF, gSDF, MOHO, and Our Method on in-the-wild Images
@@
9.4 Failure Cases Analysis
Most failures originate from the initial reconstruction stage rather than the optimization stage. As shown in Figure 14, the inpainting model occasionally introduces artifacts, causing the object being held to blend with the background and making segmentation challenging. Consequently, the input images for the large reconstruction model become incoherent and deviate significantly from real-world objects. We believe that incorporating more advanced inpainting models and leveraging hand contours in the segmentation process are promising directions for future exploration.
Figure 14: Examples of failure Cases in the OakInk dataset. Failures primarily result from artifacts introduced by the inpainting model.
We also evaluated our pipeline on the MOW dataset. The images in this dataset are of relatively low quality, frequently displaying ambiguous grasping poses and motion blur. These factors present significant challenges for both inpainting and object reconstruction. Figure 15 illustrates typical examples of failure cases observed on the MOW dataset.
Figure 15: Examples of failure cases on the MOW dataset. The last column is the synthesized 6 views from InstantMesh, it shows that motion blur and ambiguous grasping present significant challenges for object reconstruction.